A bag contains 5 red nd 3 blue balls. if 3 balls are drawn at random without replacement the probability that exactly two of three balls were red, the first ball being red is??

Grade:12

3 Answers

Rinkoo Gupta
10 years ago
Total no of ways of selecting 3 balls out of total 8 balls=8C3
Total no of selecting 2 red ball out of 5 red ball and 1 blue ball out of 3 blue balls=5C2*3C1
Hence the probability of exactly 2 red balls out of 3 balls=(5C2*3C1)/8C3
=(10)(3)/56=30/56=15/28
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty
Pushkar Aditya
10 years ago
This is correct answer.
VISHAL GAUTAM
10 years ago
Total no. of selections=5C1*7C2 Total possible selections =5C2*3C1 Hence probablity=5C2*3C1 /5C1*7C2 =2/7